# Learning Generative Interactive Environments By Trained Agent Exploration

### v1

@@ Abstract
World models are increasingly pivotal in interpreting and simulating the rules and actions of complex environments. Genie, a recent model, excels at learning from visually diverse environments but relies on costly human-collected data. We observe that their alternative method of using random agents is too limited to explore the environment. We propose to improve the model by employing reinforcement learning based agents for data generation. This approach produces diverse datasets that enhance the model’s ability to adapt and perform well across various scenarios and realistic actions within the environment. In this paper, we first release the model GenieRedux- an implementation based on Genie. Additionally, we introduce GenieRedux-G, a variant that uses the agent’s readily available actions to factor out action prediction uncertainty during validation. Our evaluation, including a replication of the Coinrun case study, shows that GenieRedux-G achieves superior visual fidelity and controllability using the trained agent exploration. The proposed approach is reproducable, scalable and adaptable to new types of environments. Our codebase is available at https://github.com/insait-institute/GenieRedux.
@@ 1 Introduction
In this work, we first aim to implement Genie’s components Bruce et al., (2024), since Genie’s official codebase is not available. The resulting model we release under the name GenieRedux. As the trained agent gives us agent actions, we use a guided variant of the model named GenieRedux-G where the next frame prediction is conditioned on agent actions rather than on predictions from the Latent Action Model. This allows us to evaluate our proposed environment exploration while we factor out any action prediction noise. Architectures are shown on Fig. 1. We show that our model performs well both on visual fidelity and controllability. We implement the Coinrun Cobbe et al., (2019) case study, proposed by Bruce et al., (2024), with both a random agent and a trained agent and show that the latter produces a model able to perform better in diverse situations in the environment. Our setup is easily reproducable and scales up when extending to different types of environments for training.
@@ 2 Methodology
All our models are trained on 64x64 resolution with sequence size of 16, with a patch size 4. For evaluation, we use a sequence size of 10. We first train the tokenizer and the LAM separately. We then train the dynamics using frame tokens and predicted actions for GenieRedux or agent actions for GenieRedux-G. The random exploration dataset is used to obtain the GenieRedux-Base and GenieRedux-G-Base baseline models. We then fine-tune the tokenizer and LAM on the trained agent dataset, and fine-tune the dynamics to create the GenieRedux-TA and GenieRedux-G-TA models. Further details are in App. A.
@@ 3 Experiments
LAM-Base - 31.72 0.89
GenieRedux-Base - 19.72 0.65
GenieRedux-G-Base 20.56 23.11 0.72
@@
In this experiment we repeat the original case study with a random agent, as advised by Bruce et al., (2024) and evaluate our implementation of the GenieRedux-Base and GenieRedux-G-Base models and their components on the Basic Test Set. We show visual fidelity results on Tab. 2. FID is not reported for some models due to slight image processing differences, with other metrics unaffected. We note that in the original case study of Genie scores are not reported. However, we compare our tokenizer’s 38.25 PSNR with the reported tokenizer’s 35.7 PSNR in their Appendix C.2. We also observe high visual fidelity for our LAM. In spite of that, we observed LAM to struggle with capturing actions. This lead to insufficient action conditioning for the dynamics of GenieRedux-Base. However, GenieRedux-G-Base demonstrates high quality visual fidelity, able to consistently respond to actions and progressing motions over time (demonstrated in App. B). Note that the evaluation of dynamics consists of predicting 10 images in the future, given a single image and the actions to perform. The prediction a single step one with 25 MaskGIT iterations.
Table 3: Visual Fidelity Evaluation of GenieRedux-G and its tokenizer, trained with random agent exploration (-Base), compared to training with trained agent exploration (-TA). Evaluation is done on Diverse Test Set.
@@
GenieRedux-G-Base 22.59 21.31 0.68 0.81
GenieRedux-G-TA 12.75 24.79 0.76 1.09
@@
In this experiment, we evaluate our models trained with the trained agent exploration, rather than the random agent - GenieRedux-TA and GenieRedux-G-TA. The evaluation set is the Basic Test Setto match the classic case study. Visual fidelity results are shown in Tab. 2. Tokenizer-TA and LAM-TA show significantly improved visual fidelity compared to the Base models. While LAM still struggles to represent most of the actions, we mark improvement as the jump action is better captured. GenieRedux-TA is negatively affected by LAM’s uncertainty, but it is able to perform the jump action consistently (see App. C). Meanwhile, GenieRedux-G-TA shows good visual quality and is consistently able to enact all environment actions and progress motions, as seen on Fig. 4 (more in App. E). All actions are demonstrated on Fig. 2.
@@
In this experiment we select our best models so far and compare trained exploration (GenieRedux-G-Base) with random agent exploration (GenieRedux-G-TA) on the various scenarios in the Diverse Test Set. Tab. 3 shows that trained agent exploration outperforms random exploration in terms of visual fidelity. Moreover, it offers a significant gain in controllability, when comparing with the $\Delta_{t}$ metric, defined in Bruce et al., (2024). This is also demonstrated on Fig. 2.
@@
We compare with Jafar Willi et al., (2024) - a concurrent with ours implementation of Genie (in JAX). We obtain and train their model as instructed. We train GenieRedux-Base with Jafar’s model parameters. Despite GenieRedux-Base not representing actions well, it shows significantly better visual fidelity metrics, achieving 17.91 PSNR (46.12 FID), compared to Jafar’s 12.66 PSNR (154.12 FID) . GenieRedux-Base does not exhibit Jafar’s artifacts or the reported problematic “hole digging” behavior (more in App. D). Moreover, we observe that Jafar lacks causality, which poses issues when predicting sequence content from input.
@@
We evaluate our best model’s controllability over varying prediction horizons on Fig. 3. As expected, predictions become more challenging further into the future. The first prediction is also difficult due to insufficient motion information - we obtain 0.4 $\Delta_{t}$ for $t=1$. To address this issue, we provide the model with 4 frames and actions, and observe an improvement from 24.79 PSNR (12.75 FID) to 25.88 PSNR (12.41 FID) on Diverse Test Set.
@@ 4 Conclusion
In this work, we revisited Bruce et al., (2024)’s approach, noting that while Genie achieves strong results, it relies on costly human data and limited random agent exploration. We address these limitations by demonstrating that RL-based exploration provides a scalable, effective alternative, enhancing the generalizability and efficiency of world models in complex environments.
@@ Appendix A Appendix: Training Setup
We train the tokenizer and the LAM on 6 A100 GPUs each for 100k iterations - 4 days. We finetune them for 50k iterations - 2 days. We train the dynamics model on 7 A100 GPUs for 100k iterations - 6 days. We finetune for 50k iterations - 3 days.
@@ Appendix C Appendix: GenieRedux-TA Qualitative Evaluation
On Fig. 6 we demonstrate that GenieRedux-TA is able to execute the jump action.
Figure 6: GenieRedux-TA Jump Demonstration. We present a few samples where on each iteration we perform a jump action. We demonstrate that every time GenieRedux-G-TA executes the jump action.

### v2

@@ Abstract
World models are increasing in importance for interpreting and simulating the rules and actions of complex environments. Genie, a recent model, excels at learning from visually diverse environments but relies on costly human-collected data. We observe that their alternative method of using random agents is too limited to explore the environment. We propose to improve the model by employing reinforcement learning based agents for data generation. This approach produces diverse datasets that enhance the model’s ability to adapt and perform well across various scenarios and realistic actions within the environment. In this paper, we first build, evaluate and release the model GenieRedux - a complete reproduction of Genie. Additionally, we introduce GenieRedux-G, a variant that uses the agent’s readily available actions to factor out action prediction uncertainty during validation. Our evaluation, including a replication of the Coinrun case study, shows that GenieRedux-G achieves superior visual fidelity and controllability using the trained agent exploration. The proposed approach is reproducable, scalable and adaptable to new types of environments. Our codebase is available at https://github.com/insait-institute/GenieRedux.
@@ 1 Introduction
In this work, we first reproduce the Genie model Bruce et al., (2024), as Genie’s official codebase is not available. The resulting model we release under the name GenieRedux. As the trained agent gives us agent actions, we use a guided variant of the model named GenieRedux-G where the next frame prediction is conditioned on agent actions rather than on predictions from the Latent Action Model. This allows us to evaluate our proposed environment exploration while we factor out any action prediction noise. Architectures are shown on Fig. 1. We show that our model performs well both on visual fidelity and controllability. We implement the Coinrun Cobbe et al., (2019) case study, proposed by Bruce et al., (2024), with both a random agent and a trained agent and show that the latter produces a model able to perform better in diverse situations in the environment. Our setup is easily reproducable and scales up when extending to different types of environments for training.
@@ 2 Methodology
All our models are trained on 64x64 resolution with sequence size of 16, with a patch size 4. For evaluation, we use a sequence size of 10. We first train the tokenizer. We then train the LAM and dynamics together, using frame tokens and predicted actions for GenieRedux or ground truth agent actions (no LAM) for GenieRedux-G. The random exploration dataset is used to obtain the GenieRedux-Base and GenieRedux-G-Base baseline models. We then fine-tune the tokenizer and LAM on the trained agent dataset, and fine-tune the dynamics to create the GenieRedux-TA and GenieRedux-G-TA models. Further details are in App. A.
@@ 3 Experiments
LAM-Base 37.01 33.97 0.92
GenieRedux-Base 21.88 25.51 0.77
GenieRedux-G-Base 18.88 33.41 0.92
@@
In this experiment we repeat the original case study with a random agent, as advised by Bruce et al., (2024) and evaluate our implementation of the GenieRedux-Base and GenieRedux-G-Base models and their components on the Basic Test Set. We show visual fidelity results on Tab. 2. We note that in the original case study of Genie scores are not reported. However, we compare our tokenizer’s 38.25 PSNR with the reported tokenizer’s 35.7 PSNR in their Appendix C.2. Our LAM is able to learn environment actions, leading to the visual fidelity of GenieRedux-Base, validating the correctness of our implementation. However, GenieRedux-G-Base demonstrates superior visual fidelity, controllability and ability to progress motions over time (demonstrated in App. B), as it avoids the uncertainty of LAM. Note that the evaluation of dynamics consists of predicting 10 images in the future, given a single image and the actions to perform. The prediction on a single step is with 25 MaskGIT iterations.
Table 3: Visual Fidelity Evaluation of GenieRedux, GenieRedux-G and their tokenizer, trained with random agent exploration (-Base), compared to training with trained agent exploration (-TA). Evaluation is done on Diverse Test Set.
@@
GenieRedux-Base 23.97 23.82 0.73 -
GenieRedux-G-Base 19.51 31.66 0.90 0.70
GenieRedux-TA 12.57 31.97 0.90 -
GenieRedux-G-TA 12.40 34.44 0.92 1.89
@@
In this experiment, we evaluate our models trained with the trained agent exploration, rather than the random agent - GenieRedux-TA and GenieRedux-G-TA. The evaluation set is the Basic Test Set to match the classic case study. Visual fidelity results are shown in Tab. 2. Tokenizer-TA shows significantly improved visual fidelity compared to the Base model. LAM-TA shows reduced visual fidelity which does not affect GenieRedux-TA, as performance is on-par with Base - a sign for a good predicted action quality. (see App. C). Meanwhile, GenieRedux-G-TA, unaffected by LAM’s uncertainty, shows significantly better visual quality and is consistently able to enact all environment actions and progress motions, as seen on Fig. 4 (more in App. E). All actions are demonstrated on Fig. 2.
@@
Here we compare all our models on the various scenarios in the Diverse Test Set. Tab. 3 shows that both trained agent exploration models outperform the random exploration models in terms of visual fidelity. Moreover, trained agent exploration offers a significant gain in controllability, represented by the $\Delta_{t}$ metric, defined in Bruce et al., (2024). This is also demonstrated with our best model GenieRedux-G-TA on Fig. 2.
@@
We compare with Jafar Willi et al., (2024) - a concurrent with ours implementation of Genie (in JAX). We obtain and train their model as instructed. We train GenieRedux-Base with Jafar’s model parameters and like them separate LAM from Dynamics in training. The latter significantly worsened GenieRedux-Base’s action representation. Despite that, GenieRedux-Base shows significantly better visual fidelity metrics, achieving 17.91 PSNR (46.12 FID), compared to Jafar’s 12.66 PSNR (154.12 FID) . GenieRedux-Base does not exhibit Jafar’s artifacts or the reported problematic “hole digging” behavior (more in App. D). Moreover, we observe that Jafar lacks causality which we find problematic.
@@
We evaluate our best model’s controllability (at 50k iterations) over varying prediction horizons on Fig. 3. As expected, predictions become more challenging further into the future. The first prediction is also difficult due to insufficient motion information - we obtain 0.4 $\Delta_{t}$ for $t=1$. To address this issue, we provide the model with 4 frames and actions (predicting 10), and observe an improvement of our best model (GenieRedux-G-TA) from 34.79 PSNR (12.75 FID) on Tab. 3 to 38.31 PSNR (12.29 FID) on Diverse Test Set.
@@ 4 Conclusion
In this work, we revisited Bruce et al., (2024)’s Genie - while achieving strong results, we note it relies on costly human data and limited random agent exploration. We address these limitations by demonstrating that RL-based exploration provides a scalable, effective alternative, enhancing the generalizability and efficiency of world models in complex environments.
@@ Appendix A Appendix: Training Setup
We train the tokenizer on 6 A100 GPUs for 100k iterations - 4 days. We finetune it on the trained exploration data for 150k iterations - 2 days. We train GenieRedux and GenieRedux-G models on 8 A100 GPUs for 150k iterations - 4 days.
@@ Appendix C Appendix: GenieRedux-TA Qualitative Evaluation
On Fig. 6 we demonstrate that GenieRedux-TA is able to execute actions and complete motion. On Fig. 7 we show that the model is capable of executing all actions of the environment.
Figure 6: GenieRedux-TA Qualitative Compatison. We present a few samples from the test set with various actions. We demonstrate that GenieRedux-G-TA performs the actions correctly.
Figure 7: GenieRedux-TA Controllability. We show predictions for all environment actions of GenieRedux-TA.